# Twists, Humps, and Pebbles: The Uneven Landscape of End-to-end Speech Recognition and Demographics

### v1

@@ A.5 Dataset Statistics
Table 4: Statistic on the sample used for our analysis (validation and test) for Mozilla Common Voice. Total number of records, average length in seconds and number of tokens as tokenized by Whisper’s pretrained tokenizer, number of male and female speakers, and gini dispersion index of the snippets-per-speaker dispersion, separately by subset.
@@
Table 5: Statistic on the sample used for our analysis (validation and test) for VoxPopuli. Total number of records, average length in seconds and number of tokens as tokenized by Whisper’s pretrained tokenizer, number of male and female speakers, and gini dispersion index of the snippets-per-speaker dispersion, separately by subset.
@@
Table 6: Statistic on the sample used for our analysis (validation and test) for Fleurs. Total number of records, average length in seconds and number of tokens as tokenized by Whisper’s pretrained tokenizer, number of male and female speakers, and gini dispersion index of the snippets-per-speaker dispersion, separately by subset. Speakers extracted automatically.
Table 4, 6, and 5 reports all collected statistics on CV, Fleurs, and VoxPopuli, respectively.
A.6 Details on Gender Probing
@@
A.7 Energy Statement
Appendix B Artifacts Statement
@@
  • Per-record annotations of segments with voice activity extracted from the voice activity detection pipeline (§A.2).
  • Per-record acoustic features in our phonetic study (§5).
  • Per-record embeddings and cluster IDs extracted on Fleurs with the speaker identification pipeline (§A.3).
@@
(a) Probes trained on untouched labels.
(b) Probes trained on shuffled labels.
Figure 9: F-M gender probing F1 Macro performance for every contextual embedding in Whisper-large-v3 within the first 10 seconds (x axis). Logistic regression probe with L2 regularization trained on standard (8(a)) and shuffled (8(b)) training labels. Orange lines indicate the average length of test segments. All CV languages.

### v2

@@ A.6 Dataset Statistics
Table 5: Statistic for CV. We used validation and test sets. Total number of records, average length in seconds, number of tokens as tokenized by Whisper’s pretrained tokenizer, number of unique speakers and, gini index of the snippets-per-speaker dispersion for male (M), female (F), and “other” (O) subgroups.
@@
Table 6: Statistic for VoxPopuli. We used validation and test sets. Total number of records, average length in seconds, number of tokens as tokenized by Whisper’s pretrained tokenizer, number of unique speakers and, gini index of the snippets-per-speaker dispersion for male (M) and female (F) subgroups.
@@
Table 7: Statistic for Fleurs. We used validation and test sets. Total number of records, average length in seconds, number of tokens as tokenized by Whisper’s pretrained tokenizer, number of unique speakers and, gini index of the snippets-per-speaker dispersion for male (M) and female (F) subgroups. Speakers are extracted automatically.
Table 5, 7, and 6 reports all collected statistics on CV, Fleurs, and VoxPopuli, respectively.
A.7 Energy Statement
Appendix B Details on the exploratory analysis
Dataset Model Feature rho p
VoxPopuli Seamless Sp. rate -0.10 0.76
Intensity -0.43 0.18
Pitch -0.83 0.00
Whisper Sp. rate -0.33 0.32
Intensity 0.23 0.50
Pitch -0.30 0.36
Fleurs Seamless Sp. rate 0.62 0.00
Intensity 0.11 0.66
Pitch -0.06 0.82
Whisper Sp. rate 0.20 0.42
Intensity -0.26 0.29
Pitch -0.53 0.02
CV Seamless Sp. rate -0.51 0.02
Intensity -0.25 0.31
Pitch 0.20 0.41
Whisper Sp. rate -0.39 0.10
Intensity -0.10 0.69
Pitch 0.38 0.11
Table 8: Correlation between Acoustic Features and Error Rate Gaps. Pearson correlation between the difference of the mean of subgroups an and $E(r_{F},r_{M})$. In bold the statistically relevant correlation ($p<0.05$).
ar ca cs de en es fi fr hu it ja nl pt ro ru sk sr sw yo
Fleurs
Intensity
Pitch
Speaking rate $\times$ $\times$
VoxPopuli
Intensity $\times$ $\times$
Pitch
Speaking rate $\times$ $\times$ $\times$
CV
Intensity $\times$ $\times$ $\times$
Pitch
Speaking rate $\times$ $\times$ $\times$ $\times$ $\times$
Table 9: Statistical differences of acoustic features between genders, separately by language and dataset. ✓indicates statistical difference, $\times$ indicates no difference (independent-sample Student’s T-test, $p<0.05$).
Figure 7: Acoustic features. Distributions of mean intensity, mean pitch, and speaking rate for Italian and Slovak on our three datasets. “Male” (green, solid line), “Female” (red, dashed line).
B.1 Acoustic Analysis
We provide further details about the acoustic analysis carried out in §5, comparing male and female speakers. Specifically, Table 9 reports the statistics of the T-tests conducted between acoustic features comparing genders across different languages and datasets, and Table 8 presents statistics on the Pearson correlation between differences in acoustic features and error rate gaps, separately for each (dataset, model) configuration.
We discuss the results of two representative languages only: Italian and Slovak. Italian exemplifies languages with comparable performance across genders, whereas Slovak is biased towards men and belongs to the set of languages with larger gender gap trends (see §4 for details). For both languages, acoustic features are, in most cases, statistically different. As shown in Figure 7, the most significant differences are found in mean pitch values, confirming the well-known differences between males and females. The distributions align with ranges suggested by the literature (Simpson, 2009) for both languages and all datasets. In contrast, differences in speaking rate are much less evident but still statistically significant according to the T-test (see Table 9). Intensity shows the most variability across all dimensions, even within gender groups, particularly noticeable in the Fleurs dataset. [^20] This variability could be attributed to the underlying recording conditions of Fleurs, details of which were not provided in the original paper. In Italian (VoxPopuli), there is no difference in intensity between the two gender groups ($p=0.356$), and similarly in Slovak (CV) ($p=0.376$). However, for both languages, fitting an OLS regression at the sentence level using acoustic features to predict sentence-level error rates ($r_{F}$, $r_{M}$) showed no significant contribution, similar to other languages (R2 with max: $0.20$ and averageσ: $0.03_{\pm 0.04}$)
B.2 Lexical Analysis
To capture lexical phenomena, we extracted from the reference transcript of each record two sets of features. Part-of-speech tags include NOUN, PROPN, VERB, ADJ, ADV, PRON, AUX, CCONJ, DET, PART. Named entities tags include F_LOC, F_ORG, F_PER, and MISC. For each language, we used the corresponding SpaCy (https://spacy.io/models) “medium” text analyzer (e.g., for English en_core_web_sm and de_core_news_sm for German) when available. For hu an ar we used models available in Stanza (Qi et al., 2020). For yo and sw we limited the analysis to POS tags due to NE parses unavailability. We used mbeukman/xlm-roberta-base-finetuned-ner-yorub and mbeukman/xlm-roberta-base-finetuned-swahili-finetuned-ner-swahili for yo and sw, respectively (Beukman and Fokam, 2023). We left sr, sk, and cs out of this analysis since no taggers were available. We normalized tags from different modes accordingly (e.g., we consider Stanza’s PERSON and SpaCy F_PER the same feature). Moreover, we computed lexical complexity as specified in Imani and Habil (2017), i.e., the ratio between the sum of [NOUN, PROPN, VERB, ADJ, ADV] and the sum of [PRON, AUX, CCONJ, DET, PART].
B.3 Probing Analysis
@@
Figure 8: Logistic probe F1S vs. $E(r_{F},r_{M})$. CV. Pearson’s correlation (purple dashed line) is -0.65.
Appendix C Release Statement
@@
  • Segment-level annotations of segments with voice activity extracted from the voice activity detection pipeline (§A.3).
  • Segment-level acoustic features from our acoustic analysis (§5).
  • Segment-level embeddings and cluster IDs extracted on Fleurs with the speaker identification pipeline (§A.4).
@@
(a) Probes trained on the original labels.
(b) Probes trained on the shuffled labels.
Figure 9: F-M gender probing F1 Macro performance for every contextual embedding in Whisper-large-v3 within the first 10 seconds (x axis). Logistic regression probe with L2 regularization trained on standard (LABEL:sub@sfig:logreg_fm_normal) and shuffled (LABEL:sub@sfig:logreg_fm_random) training labels. Orange lines indicate the average length of test segments. All CV languages.